OpenStudy (anonymous):

Find the family of ortogonal curves from the family: y^2=4(x-a)

OpenStudy (anonymous):

Please help me it's for a test. The answer options are: a) y=be^(x/2) b) y=be^(-x) c) y=be^(-x/2)

OpenStudy (anonymous):

first differentiate implicitly: 2yy'=4 for the orthogonal curve the slope will be -1/y'. Put this value in the equation: 2y(-1/y')=4 separate variables and solve for y: y'/y=1/2 or dy/y=-dx/2 Ln|y|=-x/2 + C or y= e^(-x/2 +C)=e^(-x/2)*e^b=e^(-x/2)*b so answer a)

OpenStudy (anonymous):

sry, :) c)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

you wellcome